the second angle of an architect's triangle is three times as large as the first. the third angle is 30 degress more than the first find the measure of each angle?
Answer by doukungfoo(195)

Note: the sum of the angles of any triangle will always equal 180 degrees
Let's call the first angle x
Given: The second angle is three times as large as the first
Second angle = 3x
Given: The thrid angle is 30 degrees more than the first
Third angle = x+30
Now we can create an equation to solve for the
first angle which we are calling x
(firstangle) + (second angle) + (third angle) = 180
(x) + (3x) + (x+30) = 180
x + 3x + x + 30 = 180
5x + 30 = 180
5x + 30 - 30 = 180 - 30
5x + 0 = 150
5x = 150
5x/5 = 150/5
x = 30

Now that we have found the first angle to be 30 degrees we can
use that to find the other two angles

Angle two = 3x
Angle two = 3(30)
Angle two = 90 degrees

Angle three = x+30
Angle three = 30+30
Angle three = 60 degrees

Now check answers
(firstangle) + (second angle) + (third angle) = 180
30 + 90 + 60 = 180
180 = 180